Write an equation of the line that passes through the point (3, 6) and has a slope of 4/9

Answer by nerdybill(7384)

Write an equation of the line that passes through the point (3, 6) and has a slope of 4/9
.
Plug into the "point-slope" form:
y - y1 = m(x - x1)
y - 6 = (4/9)(x - 3)
y - 6 = (4/9)x - (4/9)3
y - 6 = (4/9)x - 4/3
y = (4/9)x - 4/3 + 6
y = (4/9)x - 4/3 + 18/3
y = (4/9)x + 14/3